OpenStudy (anonymous):

GENIUSES!!!! A certain arithmetic sequence has the recursive formula an = an-1 + d. If the common difference between the terms of the sequence is -13, what term follows the term that has the value 13?

OpenStudy (anonymous):

an=an-1 +d an=13+(-13)=13-13=0

OpenStudy (anonymous):

can you explain it to me?

OpenStudy (anonymous):

recursive formula basically shows the relationship between the term before and the next term, so the two consecutive terms. What he did is basically just plug in value 13 and findthe next term using the recursive formula.

OpenStudy (anonymous):

ohhh okay, so the next term would be 0?

OpenStudy (anonymous):

Yes it would be, the common difference of -13 is also same as the value d, so you basically just have to put in numbers.

OpenStudy (anonymous):

Thanks A lot!